Macros for Mathematics in Plain TEX

A Macro Package by Russell Lyons

                                   rdlyons@indiana.edu

Overall Formatting

The output produced by the settings in the file \math.macros turns into very close to one page per printed page in most journals. In particular, the file sets \magnification=1200 for easier reading, \baselineskip=15pt since 12pt is too close together for mathematics, and \overfullrule=0pt, since one usually wants to accept small transgressions of the right margin. If one wants a different magnification, recall that one cannot reset it. Instead, either change this line in the file (it’s almost the first line), or make a copy of math.macros that has this line missing.

Font Environments

When one wants an entire box, line, or section to be in a different size font than the rest of the document, one wants all the fonts used to be changed at once. Also, scripts should change accordingly. This is what font environments do. These are completed from the ones in the TEXbook macros, manmac.tex. Recall that this is completely separate from the magnification. The font families supported are rm (roman), it (italic), sl (slant), bf (bold), tt (typewriter), ss (sans serif), sc (small caps), bfi (bold italic), scr (script), and math fonts. The sizes are 10, 9, and 8 points. To invoke them, just type \tenpoint, \ninepoint, or \eightpoint. There is also \fourteenpointbold for titles. The AMS symbols from amssym.def and amssym.tex are automatically input.

Some additional very large fonts are defined: \ig and \igg are roman at 14.4pt and 17.28pt, while \igb and \iggb are the corresponding bold roman fonts.

Footnotes

The ordinary footnote macro can produce some ugly output. A new one is \ftnote and is invoked just like \footnote, i.e., \ftnote {number}{text}. To number footnotes automatically, use instead \Fnote {text}.

Particular Formatting

In order to put a headline on page 1 and have other headlines blank, use \head {left side}{right side}. The macro \firstheader is defined as a default with \versiondate to be used with Ref.macros. Set the title in 14pt bold with \title {title} and the author in small caps with \author {author-list}. Put the abstract in 9pt with \abstract {abstract-text}. Put bottom material, on the first page presumably, with \bottomIII {2010-MSC-classification}{key-words}{all-else}.

An item such as a theorem can be begun by \proclaim item-type. text. The first argument ends with a period and a space, and the second argument ends with a blank line (or a \par).

To put in a long comment, use \comment {comment}. If you want to put in a manuscript note that will call attention to itself in \proofmode in Ref.macros and disappear otherwise, use \msnote {note}.

To include items indented by 25pt with the item number in roman font, type something such as

    \beginitems

\itemrm {item-nm} item-text \itemrm {item-nm} item-text \itemitemrm {itemitem-nm} itemitem-text \itemitemrm {itemitem-nm} itemitem-text

    \enditems

To put a bullet before every paragraph, type

\beginbullets

paragraphs

\endbullets

To put material in 8pt and narrower, type

\technical

paragraphs

\endtechnical

In a paper, one way to give bibliographic references at the end is to use the following macros: Begin the section with \beginreferences and end it with \endreferences. This will put a centered REFERENCES and define the macros \rrf and \and. The macro \rrf is not to be used by you, but by the program. It is inserted at the beginning of each paragraph. Thus, each reference should have a separate paragraph in the format

Author (year) Title-and-the-rest. This will produce output with the author in small caps and with a period after the year. The parentheses are required and will be reproduced. If the word “and” is needed for authors, the use of \and will give the word in roman, rather than small caps. The rest of the reference will be hang indented. Also, \frenchspacing is automatic, so you do not need to worry about spaces after periods. For example,

Burton, R. \and Pemantle, R. (1993) Local characteristics, entropy and limit
theorems for spanning trees and domino tilings via transfer-impedances,
{\it Ann. Probab.} {\bf 21}, 1329--1371.

will produce Burton, R. and Pemantle, R. (1993). Local characteristics, entropy and limit theorems for spanning

trees and domino tilings via transfer-impedances, Ann. Probab. 21, 1329–1371. The macro \myaddress will give my address; if you want yours, you need to replace mine in the file!

Common Math Items

The following are mostly self-explanatory. For more details, look at their definitions in the file.

\proof
\qed

\Qed This works in both mathmode and not. \AQed This is aligned for use in \eqalignno. Only \Qed should be used with Ref.macros, which redefines it appropriately.

\remark
\Remark
\example
\Example
\Question
\solution
\note
\definition
\notation

Math Symbols

The following are mostly self-explanatory. For more details, look at their definitions in the file.

\esssup
\essinf
\implies

\st This produces a nicely spaced semicolon for use in sets. \buldot This is a medium-size dot (“bullet dot”). \cbuldot This is a centered bullet dot. \OC {endpoint1, endpoint2} \CO {endpoint1, endpoint2} These allow half-open intervals without unmatched parentheses in the TEX file. \texists {existentially-quantified-variable} \all {universally-quantified-variable} These add an \enspace.

\Var

\I {indicated-set} \II {indicated-set-in-braces}

\N
\Q
\R
\Z
\E
\P

\CHI A slightly raised χ.

\smalloplus

\Pleft {event} \Psubleft {subscript}{event} \Pbig {event} \Pbigg {event} \PBig {event} \Eleft {random-variable} \Esubleft {subscript}{random-variable} \Ebig {random-variable} \Ebigg {random-variable} \EBig {random-variable} \Seq {sequence-item} \flr {real-number} \ceil {real-number}

\restrict

\hboxit {boxed-material} \vboxit {boxed-material}

Also, two commands are defined that enable you to switch two pairs of common symbols: \switchemptyset will interchange the current values of \emptyset and \varnothing, while \switchphi will interchange the current values of \phi and \varphi. Furthermore, the command \switchemptyset is automatically called since this is the common preference.